# Point-In-Context: Understanding Point Cloud via In-Context Learning

### v1

@@ Abstract
With the emergence of large-scale models trained on diverse datasets, in-context learning has emerged as a promising paradigm for multitasking, notably in natural language processing and image processing. However, its application in 3D point cloud tasks remains largely unexplored. In this work, we introduce Point-In-Context (PIC), a novel framework for 3D point cloud understanding via in-context learning. We address the technical challenge of effectively extending masked point modeling to 3D point clouds by introducing a Joint Sampling module and proposing a vanilla version of PIC called Point-In-Context-Generalist (PIC-G). PIC-G is designed as a generalist model for various 3D point cloud tasks, with inputs and outputs modeled as coordinates. In this paradigm, the challenging segmentation task is achieved by assigning label points with XYZ coordinates for each category; the final prediction is then chosen based on the label point closest to the predictions. To break the limitation by the fixed label-coordinate assignment, which has poor generalization upon novel classes, we propose two novel training strategies, In-Context Labeling and In-Context Enhancing, forming an extended version of PIC named Point-In-Context-Segmenter (PIC-S), targeting improving dynamic context labeling and model training. By utilizing dynamic in-context labels and extra in-context pairs, PIC-S achieves enhanced performance and generalization capability in and across part segmentation datasets. It is worth noting that PIC is a general framework so that other tasks or datasets can be seamlessly introduced into our PIC through a unified data format. We conduct extensive experiments to validate the versatility and adaptability of our proposed methods in handling a wide range of tasks and segmenting multi-datasets. Our PIC-S is capable of generalizing unseen datasets and performing novel part segmentation by customizing prompts. Code is available at https://github.com/fanglaosi/Point-In-Context.
Index Terms:
In-context learning, point cloud analysis, multi-task learning, part segmentation
@@ 1 Introduction
Figure 1: Illustration of in-context learning for different tasks. (a) In-context learning in NLP [1], with different text prompts for corresponding tasks: translation and sentiment analysis. (b) In-context learning in 2D vision [2], with 2D visual prompts for different tasks: segmentation and inpainting. (c) Our proposed Point-In-Context-Generalist (PIC-G) for 3D point cloud multitasking, with 3D visual prompts for different tasks: reconstruction, denoising, registration, etc. (d) Our proposed Point-In-Context-Segmenter (PIC-S) for multiple part segmentation datasets in 3D point clouds, including ShapeNetPart [3], Human3D [4], BEHAVE [5], and AKB-48 [6]. Note that our PIC-S can generalize to unseen segmentation dataset AKB-48, which is not included in the training set.
In recent years, there has been a notable emergence of large-scale models [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] boasting enormous parameters, having undergone pre-training on extensive datasets within the domains of Computer Vision (CV) and Natural Language Processing (NLP). These sophisticated models exhibit a remarkable capacity to tackle multiple tasks simultaneously, demonstrating their adaptability to new challenges. Text-to-image generation models like DALL-E [18] and language models such as GPTs [19, 1] exemplify this advancement. Nevertheless, the training process for these models demands significant computational resources. This factor renders full fine-tuning [20, 21, 22], or even adopting parameter-efficient tuning methodologies [15, 23, 24] like prompt tuning, impractical to a considerable portion of users. Thus, despite the promise they hold, the widespread adoption and practical utilization of these advanced models are hindered by resource constraints.
@@
In-context learning, originating from NLP [19, 16, 8], holds potential as the mainstream approach for efficient model adaptation and generalization. Unlike other methods that necessitate model parameter updates for new tasks, in-context learning incorporates domain-specific input-output pairs, known as in-context examples or prompts, into a query sample. In NLP, the prompt can be a machine translation pair or a sentiment analysis pair, as shown in Fig. 1 (a). This allows the model to produce optimal results without requiring any parameter updates for previous tasks. Several works [2, 13] explore in-context learning in CV. The visual prompt study [2] is the first to adopt a pre-trained neural network for filling missing patches in grid-like images, as shown in Fig. 1 (b). Other studies [13, 25] investigate the effect of visual prompts or generalization to more vision tasks. The above methods adopt Mask Image Modeling [26] (MIM) architecture for in-context task transfer. Inspired by MIM, Masked Point Modeling [27] (MPM) has been proposed recently and is widely used in different point cloud tasks [28, 29]. To our knowledge, no work has explored in-context learning for 3D point cloud understanding using the MPM framework. In this study, we propose Point-In-Context (PIC), the first in-context learning framework for 3D point cloud understanding, which is shown in Fig. 1 (c)(d).
@@
Given the absence of previous works in 3D vision, we propose a multitask benchmark leveraging commonly available datasets such as ShapeNet [30] and ShapeNetPart [3], including four common point cloud tasks: point cloud reconstruction, denoising, registration, and part segmentation. Meanwhile, we benchmark several representative baselines, including task-specific models customized for each task and multitask models equipped with a shared/pre-trained backbone with multiple task heads. Then, to tackle the multitask benchmark mentioned above, we first present the vanilla version of PIC, Point-In-Context-Generalist (PIC-G), a novel in-context framework for 3D point cloud multitasking, to explore in-context learning for 3D point clouds.
@@
A straightforward extension of the 3D MPM framework to point clouds for in-context learning may encounter two primary obstacles. Firstly, the conventional 3D MPM approach of employing position embeddings may introduce information leakage. As shown in Fig. 2 (a), previous pre-training pipelines embed positional information with the center point coordinates of all patches, even for those that are masked out (invisible). Since the patches masked out in the target should be invisible during inference, such an operation will cause information leakage, which does not satisfy the requirements in the in-context learning setting. Secondly, unlike the structured nature of 1D word embeddings or 2D images, 3D point cloud data, which are inherently unordered, present the risk of having their positional information disarrayed when partitioned into a patch sequence. In response to these issues, we propose a simple yet effective solution, termed Joint Sampling (JS) module. This technique involves recording the indices of sampled center points and employing the K-nearest neighbor strategy to simultaneously sample both the input and target. In addition, we explore two distinct baselines for PIC-G, which encompass separating inputs and targets akin to the Painter [13] strategy and concatenating inputs and targets analogous to the MAE [26] approach for reconstruction. It is worth noting that unlike the conventional focus of few-shot learning on a single specific task, our objective with PIC-G is to explore the in-context learning capabilities of models. This empowers the model to execute tasks aligned with the 3D prompt without updating model parameters, broadening its applicability and versatility within 3D point cloud understanding.
@@
For the part segmentation task, PIC-G converts each category into a point containing XYZ coordinates and forms a pre-defined label point map for segmentation tasks. However, this approach faces three obstacles in segmentation tasks. 1) Trained to fit fixed label contracts the essence of in-context learning, thus part segmentation is isolated from other tasks. 2) When multiple datasets are involved, a large set of labels needs to crowd into a limited coordinate space, making the label regression more challenging. 3) Fixed label-coordinate assignment sacrifices the possibility of accommodating novel labels when facing unseen datasets.
@@
To address large-scale and diverse part segmentation tasks within a single training, we further propose an extended version of PIC, Point-In-Context-Segmenter (PIC-S), which includes the In-Context Labeling and the In-Context Enhancing training strategies. In-Context Labeling is the key for PIC-S to generalize to other unseen datasets. It revolutionizes the approach to labeling by replacing fixed label coordinates with dynamic context-aware label points. By randomly assigning coordinates to part categories within each point cloud and aligning prompts with shared part semantics, PIC-S fosters a deeper understanding of semantic information while enhancing scalability and extensibility across various segmentation datasets. Furthermore, the In-Context Enhancing strategy provides more diverse point cloud pairs in various corruption operations, enabling the model to learn more robust mapping relationships from the prompts. Emphasizing in-context semantic information within prompts, PIC-S learns to generalize effectively across diverse datasets. During inference, PIC-S dynamically adapts the label-point assignments for each query point cloud based on a temporary label bank from prompts, ensuring robust performance even on unseen datasets and providing convenience for seamlessly introducing other datasets into our dataset.
@@
To explore stronger performance and better generalization of PIC on segmentation tasks, we create a new benchmark related to part segmentation in point clouds. The new benchmark include ShapeNetPart [3], Human3D [4], BEHAVE [5], and AKB-48 [6]. We process these datasets into a format suitable for part segmentation and integrate them into large-scale Human & Object Segmentation Datasets, consisting of approximately 140K samples.
@@
Our previous work is published in conference [31], and we make more significant contributions in this extension to make a full exploration of 3D point cloud in-context learning:
1) Within the 3D in-context learning framework, we further propose Point-In-Context-Segmenter (PIC-S). In particular, we propose In-Context Labeling and In-Context Enhancing to improve the performance and generalization capability in 3D point cloud part segmentation tasks. Furthermore, our PIC-S can seamlessly integrate additional segmented datasets without redundant label points. 2) We establish the Human & Object Segmentation benchmark, which comprises four available point cloud datasets on human and object segmentation, including ShapeNetPart, Human3D, BEHAVE, and AKB-48. Our goal is to fully evaluate the performance of models trained jointly on multiple segmentation datasets, as well as their generalization to unseen datasets. 3) We conduct extensive experiments on the Human & Object Segmentation benchmark to validate our PIC-S model. Compared to other models, our PIC-S achieves SOTA performance. Furthermore, we show that PIC-S can produce excellent results on out-of-domain part segmentation datasets, which makes our work more applicable in real-world scenarios.
@@ 2 Related Work
3D Point Cloud Analysis. With the increased large-scale 3D datasets [32, 33, 34, 35], research on 3D point cloud analysis mainly explores deep neural network architectures [36, 37]. Several pioneer works adopt the point-based methods, with Multi-Layer Perception (MLP), to directly process each point into high-dimensional embeddings, including PointNet [38] and PointNet++ [39]. Then, PointNeXt [40] improves PointNet++ via well-designed training strategies and an inverted residual bottleneck design. Meanwhile, graph-based methods [41, 42, 43, 44] leverage geometric properties. They propose different dynamic kernels with dynamic graph convolution architectures. In particular, DGCNN [41] designs EdgeConv to compute each layer’s output dynamically and achieves comparable results with point-based methods. Later, with the rise of vision transformers [45, 46], some works [47, 48, 49] adopt transformer-based approaches to build the global context between each point. Point Transformer series [47, 50, 51] explore modified transformer layers from different aspects, including a new vectorized self-attention mechanism, efficient and accuracy architecture, and extensive traverse functions. As a concurrent work, PCT [48] presents a global-local self-attention layer. In addition, several works [52, 53] integrate the vision-language models [52, 53] and 2D image [54, 55, 56, 57] features via transformer architectures for 3D point cloud analysis. Several works [58, 59] explore transformer architecture for scene-level point cloud understanding with larger receptive fields. Recently, there have also been several works exploring state space models [60, 61, 62]. In particular, PCM [61] shows great potential for 3D point cloud analysis and achieves comparable results with transformer-based models but with global receptive fields with linear complexity. However, these approaches are only designed for a single task for point cloud representation learning, so they cannot be used directly for in-context learning. Our work builds the first step in exploring in-context learning in the 3D point cloud multitasking and delves into the generalization of in-context learning in multi-dataset segmentation tasks.
@@
Masked X Modeling. By masking partial data and predicting the missing contents, Masked X Modeling acts as an effective pre-training method in both language modeling and computer vision tasks. In particular, starting from BERT [63] and GPTs [19, 1, 64], Masked Language Modeling (MLM) methods have greatly enhanced natural language processing performance by pre-training and fine-tuning diagrams. In computer vision, BEiT [65] is the first to propose matching image patches with discrete tokens via d-VAE [66] and pre-train a standard vision transformer (ViT [45]) using Masked Image Modeling (MIM). Later, MAE [26] shows that we can reconstruct the raw pixel values of masked tokens and achieve high efficiency with a high mask ratio. For 3D point clouds, several works [67, 27, 68, 69, 70, 71] design a Masked Point Modeling (MPM) pre-training framework. For example, Point-BERT [67] adopts a BERT-like architecture, while PointMAE [27] transfers the MAE-like framework for pre-training via reconstructing point cloud patches. Meanwhile, several works develop improved MPM-style frameworks via specific technologies, such as hierarchical sampling [29], cross-modal distillation [28, 72, 73, 74], etc. These methods all use standard transformer networks to process 3D point clouds and achieve competitive performance on various downstream tasks. Our approach follows a similar MPM pipeline but explores the in-context ability of point transformers and MPM. To our knowledge, this has not been investigated previously. In particular, via in-context learning, we can develop one unified model via a 3D task prompt, further widening the research scope in the 3D point cloud.
@@
Multi-Task Learning. Multi-task learning (MTL) [75] aims to learn multiple tasks jointly so that the knowledge contained in a task can be leveraged by other tasks and increase the average performance. Common methods design novel training strategies [76, 77] or model structures [78, 79, 80, 81, 82, 83] to share information and achieve mutual benefits across multiple target tasks. In 3D, several works [84, 85] rely on heavily-designed multiple task heads [86, 87] and the combination of various loss functions [88, 89]. For example, Hassani et al [84] propose an unsupervised multi-task model for jointly learning point and shape features in point clouds. To extract highly effective features, GPA-Net [87], comprising an encoder and multi-task decoders, proposes a novel graph convolution kernel termed GPAConv. The aforementioned methods achieve multi-task learning by incorporating multiple task heads. In contrast, our approach explores a new way to achieve multi-task learning by task prompting. We integrate several tasks into a unified input-output space, enabling training across various tasks within the same framework without task-specific head designs.
@@
In-Context Learning. This concept is a novel learning paradigm within large language models like GPT-3 [1], facilitating inference on novel tasks by conditioning inputs on specific input-output pairs, known as "prompts". This paradigm empowers users to customize a model’s output to align with their downstream datasets without modifying the often inaccessible internal model parameters [15, 16]. Recent advancements [7, 90] in language models underscore the effectiveness of in-context learning across diverse language tasks such as machine translation and sentiment analysis. Moreover, several studies [91, 92, 93, 11, 94, 95, 96, 97] delve into extending in-context learning to computer vision tasks. Notably, Visual Prompt [2] stands out as the pioneering pure 2D vision model, pre-trained to reconstruct masked patches within images comprised of academic figures and infographics. Subsequently, Painter [13] extends this concept by generalizing visual in-context learning to image painting across different prompts. Later, SegGPT [25] expands upon Painter by enabling generalized one-shot segmentation. Some works [98, 99, 100, 101] introduce in-context learning in diffusion-based generative models [102]. In contrast, our approach investigates the influence of 3D prompts on in-context learning within point clouds, proposing novel baselines for benchmarking 3D in-context learning efficacy. In addition, we further enhance the 3D prompt design on dense prediction tasks.
@@
In this section, we first model the in-context learning paradigm in 3D point clouds as the 2D in-context learning methods do (Sec. 3.1). Then we introduce a novel multitasking benchmark for 3D in-context inference based on ShapeNet [30] and ShapeNetPart [3] datasets (Sec. 3.2). Lastly, to further explore the in-context capability for 3D part segmentation, we establish a new multi-dataset joint training benchmark, including ShapeNetPart [3], Human3D [4], BEHAVE [5], and AKB-48 [6] (Sec. 3.3).

### v2

@@ Abstract
The rise of large-scale models has catalyzed in-context learning as a powerful approach for multitasking, particularly in natural language and image processing. However, its application to 3D point cloud tasks has been largely unexplored. In this paper, we introduce Point-In-Context (PIC), a pioneering framework for 3D point cloud understanding that leverages in-context learning with a standard transformer architecture. PIC uniquely enables the execution of multiple tasks after a single, unified training phase, eliminating the need for fine-tuning. To extend masked point modeling to 3D in-context learning, we introduce a Joint Sampling module, a simple yet effective technique that emphasizes the mapping relationship between input and target. PIC treats both inputs and targets as coordinate-based, addressing the segmentation challenge by associating label points with pre-defined XYZ coordinates for each category. However, relying on such fixed label-coordinate assignments limits the model’s ability to generalize to unseen domains. To address this limitation, we further propose two innovative training strategies: In-Context Labeling and In-Context Enhancing. These strategies are integrated into an extended version of our model, named PIC++, which enhances dynamic in-context labeling and model training. Besides its multitask capability, PIC++ demonstrates superior performance and generalization across part segmentation datasets by employing dynamic in-context labels and regular in-context pairs. Remarkably, PIC++, trained once without fine-tuning, can generalize effectively to unseen datasets and perform novel part segmentation through customized prompts. Overall, PIC is a general framework that seamlessly integrates additional tasks or datasets through a unified data format via in-context learning. Extensive experiments substantiate PIC’s versatility and adaptability in handling diverse tasks and segmenting multiple datasets simultaneously. The implementation of our framework is publicly available at https://github.com/fanglaosi/Point-In-Context.
keywords:
In-context learning, part segmentation, multi-dataset, self-supervised learning
Figure 1: Illustration of in-context learning for different tasks. a) In-context learning in NLP [gpt3], with different text prompts for corresponding tasks: translation and sentiment analysis. b) In-context learning in 2D vision [visualprompt], with 2D visual prompts for different tasks: segmentation and inpainting. c) Our proposed Point-In-Context (PIC) for 3D point cloud multitasking, with 3D visual prompts for different tasks: reconstruction, denoising, registration, etc. d) Our proposed Point-In-Context++ (PIC++) for multiple part segmentation datasets in 3D point clouds, including ShapeNetPart [shapenetpart], Human3D [human3d], BEHAVE [behave], and AKB-48 [akb48]. Note that our PIC++ can generalize to unseen segmentation dataset AKB-48, which is not included in the training set.
@@ 1 Introduction
In recent years, we have witnessed a remarkable surge in the development of large-scale models that boast an enormous number of parameters [flamingo, clip, scaling, opportunities, sam_iccv, OMGSeg, painter, wu2023open, learningtoprompt, learningtoretrieve, sic]. These models have undergone pre-training on extensive datasets across the domains of Computer Vision (CV) and Natural Language Processing (NLP), showcasing their ability to tackle a multitude of tasks simultaneously and adapt to new challenges with ease. Pioneering models such as DALL-E for text-to-image generation [dell-e] and versatile language models like GPTs [gpt, gpt3] exemplify the cutting-edge advancements in this field. Despite their remarkable capabilities, these sophisticated models face a significant barrier to adoption: the substantial computational resources required for their training process. Full fine-tuning [swintransformer, swintransformerv2, p-tuning] of these models is often prohibitive due to the high computational costs involved. Even more parameter-efficient tuning methodologies, such as prompt tuning [learningtoprompt, prefix, parameter], can be impractical for a considerable portion of users who lack access to extensive computational resources. Fortunately, In-context learning [gpt3] offers a promising method to eliminate the need for fine-tuning down-stream tasks.
@@
In-context learning, originating from NLP [gpt, learningtoretrieve, clip], holds potential as the mainstream approach for efficient model adaptation and generalization. Unlike other methods that necessitate model parameter updates for new tasks, in-context learning incorporates domain-specific input-target pairs, known as in-context examples or prompts, into a query sample. In NLP, the prompt can be a machine translation pair or a sentiment analysis pair, as shown in Fig. 1 (a). This allows the model to produce optimal results without requiring any parameter updates for previous tasks. Several works [visualprompt, painter] explore in-context learning in CV. The visual prompt study [visualprompt] is the first to adopt a trained neural network for filling missing patches in grid-like images, as shown in Fig. 1 (b). Other studies [painter, seggpt] investigate the effect of visual prompts or generalization to more vision tasks. The above methods adopt Mask Image Modeling [mae] (MIM) architecture for in-context task transfer. Inspired by MIM, Masked Point Modeling [pointmae] (MPM) has been proposed recently and is widely used in different point cloud tasks [i2pmae, pointm2ae]. To our knowledge, no work has explored in-context learning for 3D point cloud understanding using the MPM framework. In this study, we propose Point-In-Context (PIC), the first in-context learning framework for 3D point cloud understanding, realizing one training, one model, and processing multiple tasks, which is shown in Fig. 1 (c). Additionally, we propose the extended version of PIC, named Point-In-Context++ (PIC++), a unified in-context model specifically designed for multi-entity part segmentation tasks, which is shown in Fig. 1 (d).
@@
To systematically evaluate the 3D in-context learning capability, which serves as the foundation for both PIC and PIC++, we first propose a multitask benchmark leveraging commonly available datasets such as ShapeNet [shapenet] and ShapeNetPart [shapenetpart], including four common point cloud tasks: point cloud reconstruction, denoising, registration, and part segmentation. Meanwhile, we benchmark several representative baselines, including task-specific models customized for each task and multitask models equipped with a shared/pre-trained backbone with multiple task heads. Then, to tackle the multi-task benchmark mentioned above, we present the base framework fundamental in-context learning for 3D point clouds.
@@
A straightforward extension of the 3D MPM framework to point clouds for in-context learning may encounter two primary obstacles. Firstly, the conventional 3D MPM approach of employing position embeddings may introduce information leakage. As shown in Fig. 2 (a), previous pre-training pipelines embed positional information with the center point coordinates of all patches, even for those that are masked out (invisible). Since the patches masked out in the target should be invisible during inference, such an operation will cause information leakage, which does not satisfy the requirements in the in-context learning setting. Secondly, unlike the structured nature of 1D word embeddings or 2D images, 3D point cloud data, which are inherently unordered, present the risk of having their positional information disarrayed when partitioned into a patch sequence. In response to these issues, we propose a simple yet effective solution, termed the Joint Sampling (JS) module. This technique involves recording the indices of sampled center points and employing the K-nearest neighbor strategy to simultaneously sample both the input and target. In addition, we explore two distinct baselines for PIC, which encompass separating inputs and targets akin to the Painter [painter] strategy and concatenating inputs and targets analogous to the MAE [mae] approach for reconstruction. It is worth noting that, unlike the conventional focus of few-shot learning on a single specific task, our objective with PIC is to explore the in-context learning capabilities of models. This empowers the model to execute tasks aligned with the 3D prompt without updating model parameters, broadening its applicability and versatility within 3D point cloud understanding. We also adopt the Objaverse dataset [objaverse] (800K) to verify that integrating large-scale data into the training set can enhance performance on in-domain or out-of-domain datasets.
@@
For the part segmentation task, PIC converts each category into a point containing XYZ coordinates and forms a pre-defined label point map for segmentation tasks. However, this approach faces three obstacles in segmentation. 1) Trained to fit fixed label contracts, the essence of in-context learning, thus part segmentation is isolated from other tasks. 2) When multiple datasets are involved, a large set of labels needs to crowd into a limited coordinate space, making the label regression more challenging. 3) Fixed label-coordinate assignment sacrifices the possibility of accommodating novel labels when facing unseen datasets.
@@
To address large-scale and diverse part segmentation tasks within a single training, we further propose an extended version of PIC, PIC++, which includes the In-Context Labeling and the In-Context Enhancing training strategies. In-Context Labeling is the key for PIC++ to generalize to other unseen datasets. It revolutionizes the approach to labeling by replacing fixed label coordinates with dynamic context-aware label points. By randomly assigning coordinates to part categories within each point cloud and aligning prompts with shared part semantics, PIC++ fosters a deeper understanding of semantic information while enhancing scalability and extensibility across various segmentation datasets. Furthermore, the In-Context Enhancing strategy provides more diverse point cloud pairs in various corruption operations, enabling the model to learn more robust mapping relationships from the prompts. Emphasizing in-context semantic information within prompts, PIC++ learns to generalize effectively across diverse datasets. During inference, PIC++ dynamically adapts the label-point assignments for each query point cloud based on a temporary label bank from prompts, ensuring robust performance even on unseen datasets and providing convenience for seamlessly introducing other datasets into our dataset.
@@
To explore stronger performance and better generalization of PIC on segmentation tasks, we create a new benchmark related to part segmentation in point clouds. The new benchmark includes ShapeNetPart [shapenetpart], Human3D [human3d], BEHAVE [behave], and AKB-48 [akb48]. We process these datasets into a format suitable for part segmentation and integrate them into large-scale Multi-Entity In-Context Datasets, consisting of 140K samples.
@@
Our previous work is published in NeurIPS-2023 [pic](spotlight), and we make more significant contributions in this extension to make the first solid exploration of 3D point cloud in-context learning:
  • Within the 3D in-context learning framework, we further propose PIC++. In particular, we propose In-Context Labeling and In-Context Enhancing to improve performance and generalization capability in 3D point cloud part segmentation tasks. Furthermore, our PIC++ can seamlessly integrate additional segmented datasets without redundant label points.
  • We establish the Multi-Entity Segmentation benchmark, which comprises four available point cloud datasets on human and object segmentation, including ShapeNetPart, Human3D, BEHAVE, and AKB-48. Our goal is to fully evaluate the performance of models trained jointly on multiple segmentation datasets/tasks, as well as their generalization to unseen datasets/tasks.
  • We conduct extensive experiments on the multi-tasking and multi-dataset segmentation benchmark to validate our PIC++. Compared to other models, PIC++ achieves SoTA performance. Furthermore, we show that PIC++ can produce excellent results on out-of-domain part segmentation datasets or customized part segmentation tasks, which makes our work more applicable in real-world scenarios.
@@ 2 Related Work
3D Point Cloud Analysis. With the increased large-scale 3D datasets [modelnet40, rs, modelnet-o, wu2023omniobject3d], research on 3D point cloud analysis mainly explores deep neural network architectures [guo2020deep]. Several pioneer works adopt the point-based methods, with Multi-Layer Perception (MLP), to directly process each point into high-dimensional embeddings, including PointNet [pointnet] and PointNet++ [pointnet++]. Then, PointNeXt [pointnext] improves PointNet++ via well-designed training strategies and an inverted residual bottleneck design. Meanwhile, graph-based methods [dgcnn, lpdnet, hierarchicalgraph, lin2021learninggraph] leverage geometric properties. They propose different dynamic kernels with dynamic graph convolution architectures. In particular, DGCNN [dgcnn] designs EdgeConv to compute each layer’s output dynamically and achieves comparable results with point-based methods. Later, with the rise of vision transformers [vit, transformer], some works [pointtransformer, pct, pointr] adopt transformer-based approaches to build the global context between each point. Point Transformer series [pointtransformer, pointtransformerv2, pointtransformerv3] explore modified transformer layers from different aspects, including a new vectorized self-attention mechanism, an efficient and accurate architecture, and extensive traverse functions. As a concurrent work, PCT [pct] presents a global-local self-attention layer. In addition, several works integrate the vision-language models [pointclip, clip2] and 2D image [p2p, openscene, vg4d] features via transformer architectures for 3D point cloud analysis. Several works [lai2022stratified, Schult23ICRA] explore transformer architecture for scene-level point cloud understanding with larger receptive fields. Recently, there have also been several works exploring state space models [mamba, han2024mamba3d, pcm, pointmamba1]. In particular, MAMBA3D [han2024mamba3d] and PCMamba [pcm] show great potential for 3D point cloud analysis and achieve comparable results with transformer-based models but with global receptive fields with linear complexity. However, these approaches are only designed for a single task for point cloud representation learning, so they cannot be used directly for in-context learning. Our work builds the first step in exploring in-context learning in the 3D point cloud multitasking and delves into the generalization of in-context learning in multi-dataset segmentation.
@@
Masked X Modeling. By masking partial data and predicting the missing contents, Masked X Modeling acts as an effective pre-training method in both language modeling and computer vision tasks. In particular, starting from BERT [bert] and GPTs [gpt, gpt3, language], Masked Language Modeling (MLM) methods have greatly enhanced natural language processing performance by pre-training and fine-tuning diagrams. In computer vision, BEiT [beit] is the first to propose matching image patches with discrete tokens via d-VAE [dvae] and pre-train a standard vision transformer (ViT [vit]) using Masked Image Modeling (MIM). Later, MAE [mae] shows that we can reconstruct the raw pixel values of masked tokens and achieve high efficiency with a high mask ratio. For 3D point clouds, several works [pointbert, pointmae, pointrae, mae3d, acl, lee2022ap] design a Masked Point Modeling (MPM) pre-training framework. For example, Point-BERT [pointbert] adopts a BERT-like architecture, while PointMAE [pointmae] transfers the MAE-like framework for pre-training via reconstructing point cloud patches. Meanwhile, several works develop improved MPM-style frameworks via specific technologies, such as hierarchical sampling [pointm2ae], cross-modal distillation [i2pmae, pimae, recon, act], etc. These methods all use standard transformer networks to process 3D point clouds and achieve competitive performance on various downstream tasks. Our approach follows a similar MPM pipeline but explores the in-context ability of point transformers and MPM. To our knowledge, this has not been investigated previously. In particular, via in-context learning, we can develop a unified model via a 3D task prompt, further widening the research scope in the 3D point cloud.
Table 1: Comparison of ShapeNet In-Context Datasets and Multi-Entity In-Context Datasets with other common datasets. The underline represents the datasets we use.
Dataset Tasks In-context pair Objects Categories
ShapeNet [shapenet] 3D point cloud classification - 51k 55
ModelNet [modelnet40] 3D point cloud classification - 12k 40
3D-Future [3dfuture] 3D shape recognition, pose estimation, reconstruction, retrieval, Texture synthesis - 16k 34
ABO [abo] Single-view 3D reconstruction, Material prediction, Object retrieval - 8k 63
ScanObjectNN [scanobjectnn] 3D point cloud classification - 15k 15
ShapeNetPart [shapenetpart] 3D point cloud part segmentation - 17k 50
Human3D [human3d] 3D instance segmentation, 3D semantic segmentation, Multi-human body part segmentation - 57k 6
BEHAVE [behave] Humen-object interaction tracking, 3D reconstruction, Pose and shape estimation - 15k 20
AKB-48 [akb48] Object pose estimation, Object reconstruction, Object manipulation - 2k 48
ShapeNet In-Context Datasets 3D reconstruction, denoising, registration, part segmentation 217K 55
Multi-Entity In-Context Datasets Object part segmentation, Human part segmentation, Human-object interaction segmentation 140K 94
@@
Multi-Task Learning. Multi-task learning (MTL) [multitasksurvey] aims to learn multiple tasks jointly so that the knowledge contained in a task can be leveraged by other tasks and increase the average performance. Common methods design novel training strategies [hipro, mpa] or model structures [crossstich, learningtobranch, gnas, progressive, branched, controllable] to share information and achieve mutual benefits across multiple target tasks. In 3D, several works [unsupervisedmultitask, venvision3d] rely on heavily-designed multiple task heads [robust, gpanet] and the combination of various loss functions [jldnet, dnact]. For example, Hassani et al [unsupervisedmultitask] propose an unsupervised multitask model for jointly learning point and shape features in point clouds. To extract highly effective features, GPA-Net [gpanet], comprising an encoder and multitask decoders, proposes a novel graph convolution kernel termed GPAConv. The aforementioned methods achieve multitask learning by incorporating multiple task heads. In contrast, our approach explores a new way to achieve multitask learning by task prompting. We integrate several tasks into a unified input-target space, enabling training across various tasks within the same framework without task-specific head designs.
@@
In-Context Learning. This concept is a novel learning paradigm within large language models like GPT-3 [gpt3], facilitating inference on novel tasks by conditioning inputs on specific input-output pairs, known as ”prompts”. This paradigm empowers users to customize a model’s output to align with their downstream datasets without modifying the often inaccessible internal model parameters [learningtoprompt, learningtoretrieve]. Recent advancements [flamingo, explanation] in language models underscore the effectiveness of in-context learning across diverse language tasks such as machine translation and sentiment analysis. Moreover, several studies [exploring, incontextscene, personalize, sam_iccv, goodexample, instructmemore, selfprompting, lvm] delve into extending in-context learning to computer vision tasks. Notably, Visual Prompt [visualprompt] stands out as the pioneering pure 2D vision model, pre-trained to reconstruct masked patches within images comprised of academic figures and infographics. Subsequently, Painter [painter] extends this concept by generalizing visual in-context learning to image painting across different prompts. Later, SegGPT [seggpt] expands upon Painter by enabling generalized one-shot segmentation. Some works [icldiff, generative, instructcv, wang2024explore] introduce in-context learning in diffusion-based generative models [diff]. In contrast, our approach investigates the influence of 3D prompts on in-context learning within point clouds, proposing novel baselines for benchmarking 3D in-context learning efficacy. In addition, we further enhance the 3D prompt design on dense prediction tasks like segmentation.
@@
In this section, we first model the in-context learning paradigm in 3D point clouds as the 2D in-context learning methods do (Sec. 3.1). Then we introduce a novel multitasking benchmark for 3D in-context inference based on ShapeNet [shapenet] and ShapeNetPart [shapenetpart] datasets (Sec. 3.2). Lastly, to further explore the in-context capability for 3D part segmentation, we establish a new multi-dataset joint training benchmark, including ShapeNetPart [shapenetpart], Human3D [human3d], BEHAVE [behave], and AKB-48 [akb48] (Sec . 3.3). Tab. 1 shows the comparison of our two novel benchmarks with other commonly used datasets.